# Contrastive Disentangling: Fine-grained representation learning through multi-level contrastive learning without class priors

### v1

@@ 3.3 Disentanglement Loss
      Combine $z^{(1)}$, $z^{(2)}$ by Eq.(1) and $y^{(1)}$, $y^{(2)}$ by Eq.(5);
@@
      Calculate instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
      Calculate feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
      Calculate normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
      Combine all losses $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{inst}}+\mathcal{L}_{\text{feat}}-\alpha\cdot\mathcal{L}_{\text{entropy}}$;
      Update parameters of $f$, $g$, and $h$ to minimize $\mathcal{L}_{\text{total}}$;
@@ 4.1 Datasets
We conducted our experiments on four widely-used computer vision datasets: CIFAR-10, CIFAR-100 (Krizhevsky and Hinton, 2009), STL-10 (Coates et al., 2011), and ImageNet-10 (Deng et al., 2009). These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes used.
@@
For the CIFAR-100 dataset, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies. Furthermore, to ensure the rigor of our architecture evaluation, we did not use any unseen data during training, such as the unlabeled data in STL-10.
@@ 4.2 Experimental Settings
Data Augmentation. To ensure fair comparison with existing methods (Li et al., 2021), we adopted the same data augmentation strategies and used ResNet-34 as the backbone network for feature extraction. Specifically, all datasets underwent random cropping to 224x224, random horizontal flipping, color jittering, and grayscale transformation. Notably, Gaussian blur was applied only to the ImageNet-10 dataset, as upscaling images to 224x224 for smaller datasets (e.g., CIFAR-10, CIFAR-100, and STL-10) inherently introduces a similar blurring effect.
Training Details. In our experiments, the hidden dimension and the number of features were both set equal to the batch size. We employed a cosine decay schedule without restarts (Chen et al., 2020; Loshchilov and Hutter, 2016) and gradient clipping to enhance the stability of model training. The effectiveness of these optimizations is demonstrated in the ablation study presented in Table 2. Additionally, no other techniques (e.g., self-labeling) were used to boost model performance. The hyperparameter $\alpha$, which balances the diversity of feature representations, was simply set to 1. The remaining training settings were consistent with prior contrastive learning work (Li et al., 2021): a random seed of 42 was used, and training was conducted for 1000 epochs. The Adam optimizer (Kingma, 2014) was employed, with a learning rate of $3\times 10^{-4}$ and no weight decay. The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively.
Environment. The experiments were conducted on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor. We experimented with batch sizes of 128 and 256. The training time for the CIFAR-10 and CIFAR-100 datasets was approximately 48 GPU hours, while for the STL-10 and ImageNet-10 datasets, the training time was around 11 GPU hours. To ensure reproducibility and transparency, detailed logs of all experimental procedures are available at here.
Evaluation Metrics. We employed three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
@@ 4.3 Compared Methods
To rigorously assess the performance of our proposed method in various unsupervised learning scenarios, we compared it against a diverse set of representative methods, chosen for their relevance to the challenges addressed by our approach. Alongside the traditional K-means (MacQueen et al., 1967) clustering algorithm, we evaluated several fully unsupervised models—AE (Bengio et al., 2006), DeCNN (Zeiler et al., 2010), DAE (Vincent et al., 2010), VAE (Kingma, 2013), and DCGAN (Radford, 2015)—that span a spectrum of approaches from classical techniques to modern deep generative models. These methods, like our proposed Contrastive Disentangling (CD) model, do not rely on any prior class information during training, including the number of classes.
We also compared our method with unsupervised models that incorporate class number information—such as JULE (Yang et al., 2016), DEC (Xie et al., 2016), DAC (Chang et al., 2017), DCCM (Wu et al., 2019), PICA (Huang et al., 2020), and CC (Li et al., 2021)—to illustrate how our approach performs relative to methods that leverage this additional knowledge. To ensure a fair and consistent comparison, all experiments were conducted using the full set of samples, adhering to identical training and evaluation protocols.
@@ 4.4 Experimental Results
We evaluated our model at two key stages: feature extraction and feature disentanglement. Specifically, we applied K-means clustering to both the backbone network output and the final model output to assess the model’s capabilities in feature extraction and disentanglement, respectively.
Feature Extraction. Table 2 presents a comparative analysis of various models at the feature extraction stage, utilizing the NMI, ARI, and ACC metrics. Across four datasets—CIFAR-10, CIFAR-100-20, STL-10, and ImageNet-10—our Contrastive Disentanglement (CD) method, implemented with batch sizes of 128 (CD-128) and 256 (CD-256), demonstrated superior performance compared to other fully unsupervised learning models, including AE, DeCNN, DAE, VAE, and DCGAN. In comparison to the contrastive learning method CC, which utilizes prior information about the number of classes, CD consistently produced the highest NMI scores across all datasets, with notable improvements in STL-10 and ImageNet-10, while maintaining superior overall performance in most other cases.
Table 2: Feature extraction performance comparison across models. The 1st/2nd best results are indicated in bold/underlined
Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC CD-128 CD-256
@@
Feature Disentanglement. Table 3 presents a comparison of our method with other unsupervised methods that utilize number of classes information at the final output stage. The results demonstrate that the CD method exhibited excellent performance across all metrics, particularly on the STL-10 and ImageNet-10 datasets. For CIFAR-10 and CIFAR-100-20, although the CD method’s ARI and ACC metrics were slightly lower than those of the CC method, this may be attributed to the CD method not utilizing the number of classes information to constrain the final feature prediction. Notably, on the STL-10 dataset, while the PICA method achieved a slightly higher ACC (0.713) due to the use of additional unlabeled data, the CD-128 model reached a comparable ACC of 0.702 without using extra data, and the CD-256 model further improved this to 0.734.
Table 3: Final output performance comparison across models. The 1st/2nd best results are indicated in bold/underlined, and models that utilized unlabeled datasets during training for STL-10 are marked with
Dataset Metrics JULE DEC DAC DCCM PICA CC CD-128 CD-256
@@
These results provide strong evidence that the CD method can achieve excellent performance without requiring the number of classes information. It is particularly noteworthy that the CD method exhibited outstanding performance even with smaller batch sizes. For a detailed visualization of the CD-256 model’s feature representations using t-SNE, please see A.
@@ 4.5 Ablation Studies
To demonstrate the effectiveness of each component in our model, we conducted a series of ablation studies using the STL-10 and ImageNet-10 datasets with a batch size of 128. In sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness based on the framework and objective function shown in Figure 2. In sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
@@ 4.5.1 Normalized Entropy Loss and Batch Normalization
The impact of normalized entropy loss (NE) and batch normalization (BN) on model performance was initially evaluated, as shown in Table 4. The experimental results demonstrate that both normalized entropy loss and batch normalization significantly enhance model performance, with their synergistic application yielding even more substantial performance improvements. Notably, NE exhibited particularly significant improvements during the feature disentanglement phase, underscoring its pivotal role in facilitating effective feature disentanglement.
@@
To further illustrate the impact of normalized entropy loss on feature disentanglement, a visual representation of the model outputs is presented in Figure 3.
(a) Default
@@
Figure 3: Visualization of transposed model outputs with/without normalized entropy loss (NE loss). Each row corresponds to the prediction of a specific feature attribute across the entire batch. The left visualization (a) shows the output with NE loss enabled, focusing on common features, while the right visualization (b) shows the output without NE loss, capturing more unique, image-specific features.
This visualization clearly demonstrates the effectiveness of NE loss in guiding the model to capture common features across the dataset, rather than focusing on unique features specific to individual images. The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1 and further visualized in B. This visual evidence corroborates the quantitative results and emphasizes the significance of NE loss in the proposed methodology.
@@ 4.5.2 Training Tricks
Following the validation of core modules, we further evaluated the contribution of cosine decay schedule and gradient clipping training techniques, as shown in Table 5.
@@
The results show that both the cosine decay schedule and gradient clipping further enhance the model’s final performance and improve the stability of the results.
@@ 4.5.3 Feature Prediction Head
Next, we evaluated the effect of removing the feature prediction head, which directly reduces the model to an instance-level contrastive learning architecture, similar to SimCLR. This experiment was designed to verify the critical role of the disentanglement loss $\mathcal{L}_{\text{disentanglement}}$in the feature disentanglement process.
@@
As shown in Table 6, the results demonstrate that the feature prediction head significantly improves the model’s performance in the final layers (see Table 9 for backbone results). Removing this head leads to a noticeable degradation in performance. This confirms the critical contribution of the feature-level contrastive learning strategy, which enhances the model’s ability to disentangle features across different prediction heads.
@@ 4.5.4 Data Augmentation
This section verifies the necessity of the dual-view data augmentation strategy for feature disentanglement.
@@
As shown in Table 7, the results demonstrate that dual-view augmentation is crucial in the feature disentanglement stage for capturing common features in the dataset. However, in the feature extraction stage, single-view augmentation can achieve satisfactory performance (see Table 10 for details).
@@ 5 Conclusion
In this work, we proposed the Contrastive Disentangling (CD) framework, a novel approach for enhancing feature extraction and disentanglement in unsupervised learning without relying on prior knowledge of class numbers. The CD framework integrates multi-level contrastive learning with an innovative disentanglement loss function, emphasizing simplicity and core effectiveness. Experimental results demonstrate that CD consistently outperforms class-dependent methods with simple architectures (i.e., without additional structures such as self-labeling), particularly in NMI, across multiple benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. Moreover, CD exhibits robust performance in small-batch settings, effectively capturing common features across diverse data distributions. Comprehensive ablation studies corroborate the efficacy of each framework component, validating its overall effectiveness and potential for wide-scale application in unsupervised learning scenarios.
@@ 6 Acknowledgments
In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations (Van der Maaten and Hinton, 2008) of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. These visualizations are shown in Figures 4-7, where (a) represents the backbone output and (b) represents the feature predictor output for each dataset.
@@ Appendix B Visualizing Feature Prediction Heads with LIME
We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique (Ribeiro et al., 2016) to visualize the outputs of the feature prediction heads. LIME is configured with $num\_samples$=5000, meaning that for each image, LIME generates 5000 perturbed samples to create an interpretable explanation. We only visualize images where the feature prediction heads produce outputs greater than 0.75.

### v2

@@ 3.4 Total Loss
      Combine $\mathbf{z^{(1)}}$, $\mathbf{z^{(2)}}$ by Eq.(1) and $\mathbf{y^{(1)}}$, $\mathbf{y^{(2)}}$ by Eq.(5)
      // Compute Losses
@@
      Compute instance-level contrastive loss $\mathcal{L}_{\text{inst}}$ by Eq. (3), (4);
      Compute feature-level contrastive loss $\mathcal{L}_{\text{feat}}$ by Eq. (7), (8);
      Compute normalized entropy loss $\mathcal{L}_{\text{entropy}}$ by Eq. (9), (10);
      Compute total loss $\mathcal{L}_{\text{total}}$ by Eq. (12);
      // Update Model Parameters
       Update $f$, $g$, and $h$ by minimizing $\mathcal{L}_{\text{total}}$;
@@ 4.1 Datasets
We conducted experiments on four widely-used datasets: CIFAR-10, CIFAR-100 [20], STL-10 [21], and ImageNet-10 [22]. These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes.
@@
For CIFAR-100, we used 20 super-classes as ground-truth labels instead of the 100 fine-grained classes, aligning with benchmark settings in previous studies. Additionally, to guarantee the rigor of our assessment, no unseen data (e.g., the unlabeled data in STL-10) was used during training.
@@ 4.2 Experimental Settings
Data Augmentation. To ensure comparability with existing methods [4], we adopted the same data augmentation strategies, using ResNet-34 as the backbone for feature extraction. All datasets were subject to random cropping to 224x224, horizontal flipping, color jittering, and grayscale transformation. Additionally, Gaussian blur was applied to the ImageNet-10 dataset, while the upscaling of smaller datasets (CIFAR-10, CIFAR-100, and STL-10) provided a comparable blurring effect.
Training Details. In our experiments, both the hidden dimension and the number of features were set equal to the batch size. A cosine decay schedule without restarts [2, 23] and gradient clipping were employed to stabilize training, as validated in our ablation studies (Table 5). No additional techniques, such as self-labeling, were used to improve performance. The temperature parameters $\tau_{\text{inst}}$ and $\tau_{\text{feat}}$ were set to 0.5 and 1, respectively, while the balance parameter $\alpha$ for feature diversity was set to 1. Other training settings were based on prior contrastive learning works, including 1000 training epochs, a random seed of 42, and the Adam optimizer [24] with a learning rate of $3\times 10^{-4}$ and no weight decay.
Environment. The experiments were performed on an NVIDIA RTX 3090 GPU (24GB) and an 18 vCPU AMD EPYC 9754 128-Core Processor. Batch sizes of 128 and 256 were used. Training time for CIFAR-10 and CIFAR-100 was approximately 48 GPU hours, while for STL-10 and ImageNet-10, it was around 11 GPU hours. Detailed logs of all experimental procedures are available here.
Evaluation Metrics. We used three clustering metrics — Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Accuracy (ACC) — to evaluate the feature extraction and disentanglement capabilities of the model.
@@ 4.3 Compared Methods
We compared CD against a diverse set of representative unsupervised models, selected based on their relevance to the challenges addressed by our approach. In addition to the traditional K-means clustering algorithm [25], we assessed fully unsupervised models such as AE [7], DeCNN [26], DAE [27], VAE [8], and DCGAN [28], all of which do not rely on class information during training, like our CD model.
Furthermore, we benchmarked CD against unsupervised models that incorporate class number information, including JULE [29], DEC [30], DAC [31], DCCM [32], PICA [33], and CC [4]. All experiments were conducted using the full dataset and adhered to identical training and evaluation protocols to ensure fairness.
@@ 4.4 Experimental Results
We evaluated CD at two key stages: feature extraction and feature disentanglement. K-means clustering was applied to both the backbone output and the final model output to assess performance.
Backbone. Table 2 presents a comparative analysis at the feature extraction stage using NMI, ARI, and ACC metrics. Across all datasets — CIFAR-10, CIFAR-100, STL-10, and ImageNet-10 — both variants of CD (CD-128 and CD-256) outperformed fully unsupervised methods such as AE, DeCNN, DAE, VAE, and DCGAN. Notably, CD achieved the highest NMI scores on all datasets, with significant improvements on STL-10 and ImageNet-10, and superior overall performance compared to the class-dependent model CC.
Table 2: Feature extraction performance comparison across models. The 1st/2nd best results are indicated in bold/underlined. Models that utilized class number information during training are marked with *.
Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC* CD-128 CD-256
@@
Final output. Table 3 compares CD to models that use class number information. CD demonstrated strong performance across all metrics, particularly on more complex datasets like STL-10 and ImageNet-10. While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is expected as CD does not use class number information. Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data.
Table 3: Final output performance comparison across models. The 1st/2nd best results are indicated in bold/underlined. Models that utilized class number information during training are marked with *, while those that utilized unlabeled datasets for STL-10 are marked with .
Dataset Metrics JULE* DEC* DAC* DCCM* PICA* CC* CD-128 CD-256
@@
These results strongly support the effectiveness of CD in achieving high performance without relying on class priors. Furthermore, CD demonstrated excellent performance even with smaller batch sizes. For a t-SNE visualization of the CD-256 model’s feature representations, refer to A.
@@ 4.5 Ablation Studies
We conducted ablation studies on STL-10 and ImageNet-10 with a batch size of 128 to assess the contributions of different components of our model. In Sections 4.5.1 and 4.5.2, we incrementally added modules to validate their effectiveness, while in Sections 4.5.3 and 4.5.4, we removed certain modules to verify their necessity.
@@ 4.5.1 Normalized Entropy Loss and Batch Normalization
Table 4 demonstrates the impact of the normalized entropy loss (NE) and batch normalization (BN) on performance. Both components significantly improved results, with further enhancements observed when combined.
@@
To illustrate the effect of NE loss on feature disentanglement, Figure 3 visualizes the model’s outputs with and without NE loss. This visualization highlights the importance of NE loss in guiding the model to capture common features across the dataset.
(a) NE loss
@@
Figure 3: Visualization of transposed model outputs with/without normalized entropy loss (NE loss). The left figure (a) shows outputs with NE loss, emphasizing common features, while the right figure (b) shows outputs without NE loss, capturing more image-specific features.
The rows in Figure 3 represent the feature attributes learned by our model, as depicted in the feature-level part of Figure 1. Further visualization of these attributes using LIME is available in B.
@@ 4.5.2 Training Tricks
Next, we evaluated the contribution of cosine decay scheduling and gradient clipping (Table 5). The results show that both techniques enhanced performance and stabilized training.
@@ 4.5.3 Feature Prediction Head
We evaluated the effect of removing the feature prediction head, which reduces the model to an instance-level contrastive learning framework similar to SimCLR. As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation, highlighting the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results).
In comparison to SimCLR under identical training settings, our model demonstrates significant improvements: on the STL-10 dataset, NMI increases by 13.4%, ARI by 23.1%, and ACC by 10.9%; on ImageNet-10, NMI improves by 26.8%, ARI by 110.2%, and ACC by 53.1%. These results present that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, leading to improved clustering and representation quality in the final output.
@@ 4.5.4 Data Augmentation
Finally, we examined the necessity of the dual-view data augmentation strategy (Table 7). The results indicate that dual-view augmentation is critical for feature disentanglement, whereas single-view augmentation can still achieve satisfactory performance during feature extraction (see Table 10 for details).
@@ 5 Conclusion
In this work, we propose a novel framework that eliminates the reliance on class number assumptions, enabling the model to operate without predefined task constraints. By integrating multi-level contrastive learning with a disentanglement loss, our framework captures more fine-grained features from the data. Our experiments reveal that CD achieves comparable performance to class-dependent methods on smaller datasets (CIFAR-10, CIFAR-100), and outperforms them on more complex datasets (STL-10, ImageNet-10), despite its simple architecture. Ablation studies further confirm the effectiveness of each component. Moreover, our framework ensures that the final model outputs are enriched with semantic information, enhancing the interpretability of the learned features.
@@ 6 Acknowledgments
In this appendix, we present the t-SNE (t-Distributed Stochastic Neighbor Embedding) visualizations [34] of the outputs both the backbone feature extraction layer and the feature predictor of our model, across all datasets: CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. These visualizations are shown in Figures 4-7, where (a) represents the backbone output and (b) represents the feature predictor output for each dataset.
@@ Appendix B Visualizing Feature Prediction Heads with LIME
We utilize the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize the outputs of the feature prediction heads. LIME is configured with $num\_samples$=5000, meaning that for each image, LIME generates 5000 perturbed samples to create an interpretable explanation. We only visualize images where the feature prediction heads produce outputs greater than 0.75.
@@ Appendix C Ablation Studies on Backbone Feature Extraction
It can be observed that when evaluating the model’s ability as a pretext model, or more specifically, the feature extraction capability of the backbone, the scheduler, gradient clipping, and dual-view data augmentation are not essential. However, if we aim for the final output of the model to be meaningful, rather than focusing solely on intermediate states, these components become crucial. You may refer to Table 5 - Table 7, especially Table 6, where our model demonstrates significantly higher semantic expressiveness in the final layer.